OpenStudy (hlilly2413):

What volume (in mL) of a 50.0% (wt.) NaOH solution, that has a density of 1.53 g/mL, would you need to prepare 2.00 L of a solution that has a concentration of .200 moles of NaOH per liter. (50.0% wt. NaOH = 50.0 g of NaOH per 100 grams of solution).

OpenStudy (photon336):

@empty @rvc what do you guys think

OpenStudy (photon336):

Here's what I'm thinking. we take the concentration of NaOH and we multiply by 2.00L to find the number of moles/grams of NaOH 2.0L*(.20moL/L) = 0.40 mol of NaOH*40.0g/mol = 16grams of NaOH that's how much we I was wondering if we had say 16 grams of NaOH, to make this 50% we would also need 16 grams of solvent and we convert to ml by using the density. 50% weight would be 16 grams/(16 grams +16 grams of solution) 16/32 = 50% 1.52 g/mL = 16g/x 24.32 mL of solution

rvc (rvc):

@Empty help!

OpenStudy (abmon98):

50 wt% by weight 100 g of solution contains 50 g of NaOH. Density(g/ml)=Mass(g)/Volume(ml) Volume of Solution required to prepare-2 L Concentration- 0.200 mol/L We can Calculate the number of moles of NaOH from the given volume and concentration using the formula Number of moles(mol)=Concentration(mol/L)*Volume(L)=0.200*2=0.4 mol. We then calculate the mass of NaOH Number of moles(mol)=Mass(g)/Molar Mass(g/mol) Rearrange the formula in order to calculate the mass of your solute. Mass(g)=Molar Mass(g/mol)*Number of moles(mol) Mass(g)=40*0.4=16 g Mass of solution=16*2=32g (Since the wt% is 50) Mass of solution(g)/Density(g/ml)=32/1.53=20.9 ml

OpenStudy (photon336):

@Abmon98 interesting 50% weight Thanks for the clarification (16grams of NaOH)/(32 grams of solution total) to make it 50% 32g = 1.53g/mL ~ 32/1.53 I see what you did.